# Geolocation with Real Human Gameplay Data: A Large-Scale Dataset and Human-Like Reasoning Framework

### v1

@@ 5.3. Overall Performance Evaluation of GeoCoT
Table 3. Accuracy of different models on geolocation tasks at various scales.
@@
5.4. Ground Truth-Based Evaluation of GeoCoT Reasoning
Beyond evaluating overall task performance, we focus on analyzing the reasoning process of GeoCoT, which emulates a human-like reasoning approach. To establish a reference for this evaluation, three gaming enthusiasts collaboratively constructed reasoning processes for the same 500 cases based on geo-tagged locations, which we designated as the ground truth (GT). These GT annotations serve as a benchmark within our evaluation framework, GeoEval. The evaluation process utilizes GPT-based assessment through GPTScore (Fu et al., 2023) and prompt-based scoring across three custom-defined aspects.
The first dimension is the completeness of feature extraction (CE), which evaluates whether all key clues provided in the GT are comprehensively covered and accurately described in the reasoning process. Comprehensive feature extraction ensures that reasoning outcomes are based on sufficient factual evidence, thereby enhancing their reliability and accuracy. The second dimension is the accuracy of feature extraction (AE), which measures whether the identified and described attributes or characteristics of the key information in the GT are correct. Misidentified features can lead to reasoning outcomes that deviate from the facts, reducing the credibility of the results. The third dimension is the accuracy of reasoning and cue correspondence (AC), which assesses whether the reasoning process derives reasonable conclusions based on the extracted cues and maintains consistency with the reasoning logic presented in the GT. Incorrect correspondence between cues and conclusions can result in outcomes that deviate from reality. The final dimension is the logical coherence of reasoning (LC), which evaluates the consistency, logical flow, and adherence to common sense within the reasoning chain. Logical errors compromise the reliability of the reasoning process and hinder the model’s ability to arrive at accurate conclusions.
Table 4. Evaluation of GeoCoT’s reasoning process using ground truth-based metrics within the GeoEval framework.
@@ 5.5. Intrinsic Evaluation of GeoCoT Reasoning
The ground truth-based evaluation compares GeoCoT’s reasoning to human-crafted processes, assessing its alignment with established reasoning patterns. Moreover, to gain a more holistic understanding, we also evaluate the reasoning process from an intrinsic perspective, examining its logical consistency, coherence, and robustness without relying on external references. This intrinsic evaluation is conducted through human evaluation, as it requires multimodal assessment and certain hallucinations are difficult for VLMs to identify effectively.
Following previous work on assessing hallucinations in terms of objects, attributes, and relationships (Li et al., 2023a; Sun et al., 2023), we evaluate the quality of synthetic data across three key dimensions: (1) Object Hallucination(OH): assesses whether the synthetic data includes objects that do not exist in the image. Object Hallucination evaluates the extent to which synthetic data introduces fictional elements. (2) Fact Hallucination(FH) measures the accuracy of factual information within the synthetic data. Fact Hallucination occurs when the synthetic data contains facts, figures, or other information that is incorrect or not supported by the original data. (3) Attribution Hallucination(AH) evaluates whether the synthetic data incorrectly attributes properties, characteristics, or relations to entities or objects. To evaluate these dimensions, we invited 2 human annotators with professional backgrounds in geographic reasoning and data validation to assess GPT-4o, GeoReasoner, and our proposed GeoCoT model. These three baselines provide textual reasoning processes across 1,500 evaluated cases. The results, shown in Table 5, indicate the number of errors in each dimension, demonstrating that GeoCoT significantly reduces hallucination errors compared to the other models.
@@
Table 6. Performance comparison of GeoCoT and state-of-the-art geolocation models on traditional benchmarks.
@@ 5.7. Generalizability Evaluation
We present the performance results in Table 6. We observe that state-of-the-art geolocation models, such as GeoCLIP, achieve performance on traditional geolocation tasks that surpass GPT-4o and are almost comparable to our model, GeoCoT. However, this result is inconsistent with the performance shown in Table 2, where GeoCLIP significantly underperforms GPT-4o on fine-grained city and country-level geolocation tasks. This indicates that these baseline models are overfitting on specific datasets and highlights the importance of comprehensive evaluation across different datasets. In contrast, our model consistently outperforms traditional methods across different granularity levels and datasets.
@@ 6. Conclusion
Appendix A Detail of GeoCoT
We present the detailed prompt of our GeoCoT process below:
$\bullet$ Question1: Are there prominent natural features, such as specific types of vegetation, landforms (e.g., mountains, hills, plains), or soil characteristics, that provide clues about the geographicalregion? $\bullet$ Question2: Are there any culturally, historically, or architecturally significant landmarks, buildings, or structures, or are there any inscriptions or signs in a specific language or script that could help determine the country or region? $\bullet$ Question3: Are there distinctive road-related features, such as traffic direction (e.g., left-hand or right-hand driving), specific types of bollards, unique utility pole designs, or license platecolors and styles, which countries are known to have these characteristics? $\bullet$ Question4: Are there observable urban or rural markers (e.g., street signs, fire hydrants guideposts) , or other infrastructure elements, that can provide more specific information about the country or city? $\bullet$ Question5: Are there identifiable patterns in sidewalks (e.g., tile shapes, colors, or arrangements), clothing styles worn by people, or other culturally specific details that can help narrow down the city or area?
Let’s think step by step. Based on the question I provided, locate the location of the picture as accurately as possible. Identify the continent, country, and city, and summarize it into a paragraph. For example: the presence of tropical rainforests, palm trees, and red soil indicates a tropical climate… Signs in Thai, right-side traffic, and traditional Thai architecture further suggest it is in Thailand… Combining these clues, this image was likely taken in a city in Bangkok, Thailand, Asia.
Here, cyan highlights potential clues within the image to help the model infer geographic locations. Green defines the geographic scope inferred from the clues, such as a region, country, or city. Orange provides detailed descriptions of the cyan clues, enhancing the model’s understanding. Red specifies the expected output format, including city, country, and continent.
Appendix B Human Annotation Example
Below we show an example of human annotated ground truth to demonstrate the annotation process, criteria, and the reasoning behind the annotations, where clues are shown in blue, correct predictions in green.
The image shows a rural residential area with dense trees and expansive green lawns. The terrain is flat, and the soil is reddish-brown, which matches the temperate climate of central Europe, particularly rural areas of France. The architectural style of the house is distinctive: a red-tiled sloped roof, yellow walls, and solar panels, reflecting the region’s focus on renewable energy, a common feature in French countryside homes. The red mailbox at the gate is a hallmark of rural French residences. The design of the fences and modern gates aligns with typical styles in the French countryside. The house design and surrounding natural environment suggest a rural European region. Based on the architectural style, natural landscape, and street elements, the image was most likely taken in Aumont, France, Europe.

### v2

@@ 5.3. Overall Performance Evaluation of GeoCoT
Table 3. Accuracy of different models on geolocation tasks at various scales. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
@@
5.4. GeoEval: Reference-Based Evaluation of GeoCoT Reasoning
Beyond evaluating overall task performance, we focus on analyzing the reasoning process of GeoCoT, which emulates a human-like reasoning approach. To establish a reference for this evaluation, three gaming enthusiasts collaboratively constructed reasoning processes for the same 500 cases based on geo-tagged locations. We designated these as the reasoning ground truth (a human-annotated example can be found in Appendix C). These GT annotations serve as a benchmark within our evaluation framework, GeoEval. The evaluation process utilizes (1) GPT-based assessment through GPTScore (Fu et al., 2023) and (2) prompt-based scoring.
Our prompt-based scoring includes four dimensions ranging from 0-5, and the detailed prompts can be found in Github. The first dimension is the completeness of feature extraction (CE), which evaluates whether all key clues provided in the GT are comprehensively covered and accurately described in the reasoning process. Comprehensive feature extraction ensures that reasoning outcomes are based on sufficient factual evidence, thereby enhancing their reliability and accuracy. The second dimension is the accuracy of feature extraction (AE), which measures whether the identified and described attributes or characteristics of the key information in the GT are correct. Misidentified features can lead to reasoning outcomes that deviate from the facts, reducing the credibility of the results. The third dimension is the accuracy of reasoning and cue correspondence (AC), which assesses whether the reasoning process derives reasonable conclusions based on the extracted cues and maintains consistency with the reasoning logic presented in the GT. Incorrect correspondence between cues and conclusions can result in outcomes that deviate from reality. The final dimension is the logical coherence of reasoning (LC), which evaluates the consistency, logical flow, and adherence to common sense within the reasoning chain. Logical errors compromise the reliability of the reasoning process and hinder the model’s ability to arrive at accurate conclusions.
Table 4. Evaluation of GeoCoT’s reasoning process using ground truth-based metrics within the GeoEval framework. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
@@ 5.5. Intrinsic Evaluation of GeoCoT Reasoning
We begin with a ground truth-based evaluation, comparing GeoCoT’s reasoning to human-authored processes to assess its alignment with established reasoning patterns. To complement this, we conduct an intrinsic evaluation focused on hallucination errors—assessing logical consistency, coherence, and robustness without relying on external references. Given the need for multimodal judgment, this evaluation is performed by human annotators.
Following previous work on assessing hallucinations in terms of objects, attributes, and relationships (Li et al., 2023a; Sun et al., 2023), we evaluate the quality of synthetic data across three key dimensions: (1) Object Hallucination (OH): assesses whether the synthetic data includes objects that do not exist in the image. Object Hallucination evaluates the extent to which synthetic data introduces fictional elements. (2) Fact Hallucination (FH) measures the accuracy of factual information within the synthetic data. Fact Hallucination occurs when the synthetic data contains facts, figures, or other information that is incorrect or not supported by the original data. (3) Attribution Hallucination (AH) evaluates whether the synthetic data incorrectly attributes properties, characteristics, or relations to entities or objects. To quantify hallucinations, each detected error is counted as one instance in the corresponding dimension. To evaluate these dimensions, we invited 2 human annotators with professional backgrounds in geographic reasoning and data validation to assess GPT-4o, GeoReasoner, and our proposed GeoCoT model. These three baselines provide textual reasoning processes across 1,500 evaluated cases. The results, shown in Table 5, indicate the number of errors in each dimension, demonstrating that GeoCoT significantly reduces hallucination errors compared to the other models. The inter-annotator agreement, measured by Cohen’s Kappa, is 0.82 for OH, 0.79 for FH, and 0.85 for AH, indicating substantial agreement across all dimensions. The correlation between the human evaluation scores and the automatic evaluation metrics in Table 2 is –0.99 (p ¡ 0.01), demonstrating a strong inverse relationship: as hallucination errors decrease, overall geolocation performance improves significantly.
@@
Table 6. Performance comparison of GeoCoT and state-of-the-art geolocation models on traditional benchmarks. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
@@ 5.7. Generalizability Evaluation
We present the performance results in Table 6. We observe that state-of-the-art geolocation models, such as GeoCLIP, perform well on traditional geolocation tasks, surpassing GPT-4o and coming close to our model, GeoCoT. However, this is in contrast to the results shown in Table 2, where GeoCLIP significantly underperforms GPT-4o on fine-grained city- and country-level geolocation tasks. This discrepancy suggests that these baseline models may be overfitting to the specific datasets they were trained on, lacking the generalization ability required for more diverse or fine-grained geolocation challenges. In contrast, our model consistently outperforms traditional methods across different granularity levels and datasets without any training, and thus does not suffer from overfitting.